[figure: 9.7 cm]

MAGNETIC BEHAVIOUR AND ELECTRONIC STRUCTURE OF THE COMPOUNDS R_{2}PdSi_{3} (R = Ce, Nd, Tb-Er)

A. Szytuła^{1}, M. Hofmann^{2}, B. Penc^{1}, M. Ślaski^{3}, Subham Majumdar^{4},

E.V. Sampathkumaran^{4}, and A. Zygmunt^{5}

1 Institute of Physics, Jagiellonian University, Reymonta 4, 30-059 Krakow, Poland 2 Berlin Neutron Scattering Centre, Hahn-Meitner Institute, Berlin-Wannsee, Germany 3 School of Physics and Space Research, University of Birmingham, Birmingham, UK 4 Tata Institute of Fundamental Research, Homi Bhabha Rd., Mumbai 400005, India

5 Institute of Low Temperatures and Structural Research PAS, Wrocław, Poland The magnetic behaviour of the ternary intermetallics, R_{2}PdSi_{3} (R=Ce,Nd,Tb-Er), have been investigated by magnetization, electrical resistivity (ρ), neutron diffraction and XPS photoemission measurements. The neutron diffraction data reveal that these compounds crystallize in an AlB_{2}-type hexagonal structure and that these alloys exhibit different magnetic structures below respective magnetic transition temperatures. For Ce_{2}PdSi_{3} the Ce moments order in a sinusoidally modulated antiferromagnetic structure below 2.5 K, without extending to the entire crystal, apparently exhibiting cluster magnetism. The Nd magnetic moments in Nd_{2}PdSi_{3} exhibit ferromagnetic ordering below 17 K. In Tb_{2}PdSi_{3}, a complex situation is observed: in the temperature range 1.5 to (T_{N}=)25 K, there is a coexistence of the ferromagnetic spiral (about 75 % of magnetic moment) and spin-glass (about 25 %). For Dy, Ho and Er alloys, below T_{N}(7.5, 8 and 7 K respectively), sinusoidaly modulated structures are observed and there are upturns in ρ below T_{N} indicative of the opening up of magnetic Brillouin-zone boundary gaps in these compounds. While for Ce, the magnetic moment lies in the basal plane, for other compounds, it is parallel to the c-axis. There are distinct features in the magnetization and data near magnetic transition temperatures. The valence band and the core level photoemissions of 4d of rare earth metal, 3d Pd, 2s and 2p Si have been investigated.

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds Presentation mode : poster Corresponding author : Prof. dr hab. Andrzej Szytula Address for correspondence : Instytut Fizyki UJ 30-059 Krakow, Reymonta 4 tel (12) 632 48 88 / 5546 fax: (12) 633 70 86 Email address : szytula@if.uj.edu.pl